makaylawheeler:

Tom invests $10,000 in a savings account that offers 3.5 percent interest, compounded continuously.

Hero:

@makaylawheeler, are you familiar with the formula for continuous compound?

callielovesyhuu:

.The continuously compounded interest has this formula: A = Pe^rt

callielovesyhuu:

A = amount P = Principal e = mathematical constant e = 2.718 r = rate of interest t = time

KittyNoir:

Makayla is that you???

callielovesyhuu:

Principal = 10,000 ; rate = 3.5% ; t = 10 years. A = 10,000 * (2.718)^(0.035*10) A = 10,000 * (2.718)^0.35 A = 10,000 * 0.9513 A = 9,513 In 10 years Tom will have earned $9,513 in interest

Owen:

$9,513